Two point charges are placed as follows: charge q1 = -1.50 nC is at y = +6.00 m and charge

q2 = +3.20 nC is at the origin. What is the total force (magnitude and direction) exerted by

these two charges on a negative point charge q3 = -5.00 nC located at (2.00 m, -4.00 m)?



1
Expert's answer
2022-03-01T17:33:18-0500

"F=kq_3(\frac{q_1}{r_1^2}+\frac{q_2}{r_2^2})=8.3~nN,"

"\alpha=32°."